Factor 3m^2 – 75

this is my last one....thanks

Answer by madhan_math(34)

Factor 3m^2 – 75
Ans :
(It is simple)
3m^2-75=0
3m^2=75
m^2=75/3 = 25
i.e., m^2=25
Therefore m=5(twice)

Thus the factor is m=5,5